OpenStudy (anonymous):

Find the slope of the tangent line to the curve sqrt(3x+3y)+sqrt(2xy)=11.5 at point (4,5)

OpenStudy (tkhunny):

I like to do one thing on such a problem that many others do not emphasize. Let's check to see if the point (4,5) is actually ON the curve. Can you do that?

OpenStudy (anonymous):

yea 1 sec

OpenStudy (anonymous):

11.520708

OpenStudy (tkhunny):

Close enough? I suppose. Now what? We need a derivative, dy/dx, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but idk how to go from here can you help me this has to be submitted by 11 tonight and i am stuck

OpenStudy (tkhunny):

This is a nice Chain Rule, Implicit Derivative exercise. Go ahead and tackle the first piece. What do you get? If y = f(x), then (d/dx)(sqrt(3x+3y)) = ??

OpenStudy (anonymous):

This is a implicit derivative.

OpenStudy (tkhunny):

Excellent. What do you get for that first piece?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

(1/2)3x+3y^(-1/2)?

OpenStudy (anonymous):

1/2[(3x+3y)^-1/2](3+3y')

OpenStudy (tkhunny):

Kind of. It's pretty hard to read when written like that. You should get: \[\frac{ 3 + 3(dy/dx) }{ 2\sqrt{3x + 3y} }\] You may like it better written like this: \[\frac{ 3 + 3y' }{ 2\sqrt{3x + 3y} }\] Is that what you managed? I couldn't quite tell.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ohh wait i got the answer! thanks for the help!

OpenStudy (tkhunny):

That's only one piece. The second chunk requires Chain Rule, Implicit, and Product Rule. Can you get that one?

OpenStudy (anonymous):

i mean i was able to find solve the final answer

OpenStudy (tkhunny):

Alas. It's so much nicer when you share your work. :-)